OpenStudy (anonymous):

So do i add the 2 exponents? (14^2)^6

OpenStudy (ledah):

depends... what are you trying to do? what does the problem say?

OpenStudy (anonymous):

not in this case when you multiply you add exponents so \(b^m\times b^n=b^{m+n}\) but when you raise to a power you MULTIPLY

OpenStudy (anonymous):

. Simplify the product using the rules for exponents

OpenStudy (anonymous):

the exponents, so \((b^m)^n=b^{m\times n}\)

OpenStudy (anonymous):

in this case you would multiply the exponents not add

OpenStudy (anonymous):

A. 14^2 × 6 B. 14^8 C. 14^12 D. 14^4